# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Designing efficient algorithms for multi-agent reinforcement learning (MARL) is fundamentally challenging due to the fact that the size of the joint state and action spaces are exponentially large in the number of agents. These difficulties are exacerbated when balancing sequential global decision-making with local agent interactions. In this work, we propose a new algorithm SUBSAMPLE-MFQ (Subsample-Mean-Field-Q-learning) and a decentralized randomized policy for a system with $n$ agents. For $k\leq n$, our algorithm system learns a policy for the system in time polynomial in $k$. We show that this learned policy converges to the optimal policy in the order of $\widetilde{O}(1/\sqrt{k})$ as the number of subsampled agents $k$ increases. We validate our method empirically on Gaussian squeeze and global exploration settings.
@@ 1 Introduction
Reinforcement Learning (RL) has become a popular learning framework to solve sequential decision making problems in unknown environments, and has achieved tremendous success in a wide array of domains such as playing the game of Go (Silver et al., 2016), robotic control (Kober et al., 2013), and autonomous driving (Kiran et al., 2022; Lin et al., 2023). A critical feature of most real-world systems is their uncertain nature, and consequently RL has emerged as a powerful tool for learning optimal policies for multi-agent systems to operate in unknown environments (Kim & Giannakis, 2017; Zhang et al., 2021; Lin et al., 2024; Anand & Qu, 2024). While the early literature on RL predominantly focused on the single-agent setting, multi-agent reinforcement learning (MARL) has also recently achieved impressive successes in a broad range of areas, such as coordination of robotic swarms (Preiss et al., 2017), self-driving vehicles (DeWeese & Qu, 2024), real-time bidding (Jin et al., 2018), ride-sharing (Li et al., 2019), and stochastic games (Jin et al., 2020).
@@
Despite growing interest in multi-agent RL (MARL), extending RL to multi-agent settings poses significant computational challenges due to the curse of dimensionality (Sayin et al., 2021). Even if the individual agents’ state or action spaces are small, the global state space or action space can take values from a set with size that is exponentially large as a function of the number of agents. For example, even model-free RL algorithms such as temporal difference (TD) learning (Sutton et al., 1999) or tabular $Q$-learning require computing and storing a $Q$-function (Bertsekas & Tsitsiklis, 1996) that is as large as the state-action space. Unfortunately, in MARL, the joint state-action space is exponentially large in the number of agents. In the case where the system’s rewards are not discounted, reinforcement learning on multi-agent systems is provably NP-hard (Blondel & Tsitsiklis, 2000), and such scalability issues have been observed in the literature in a variety of settings (Guestrin et al., 2003; Papadimitriou & Tsitsiklis, 1999; Littman, 1994). Independent $Q$-learning (Tan, 1997) seeks to overcome these scalability challenges by considering other agents as a part of the environment; however, this often fails to capture a key feature of MARL, which is inter-agent interactions.
@@
Even in the fully cooperative regime, MARL is fundamentally difficult, since agents in the real-world not only interact with the environment but also with each other (Shapley, 1953). An exciting line of work that addresses this intractability is mean-field MARL (Lasry & Lions, 2007; Yang et al., 2018; Gu et al., 2021, 2022b, 2022a; Hu et al., 2023). The mean-field approach assumes that all the agents are homogeneous in their state and action spaces, enabling their interactions to be approximated by a two-agent setting: here, each agent interacts with a representative “mean agent” which is the empirical distribution of states of all other agents. Under these homogeneity assumptions, mean-field MARL allows learning of optimal policies with a sample complexity that is polynomial in the number of agents. However, when the number of homogeneous agents is large, storing a polynomially-large $Q$ table (where the polynomial’s degree depends on the size of the state space for a single agent) can still be infeasible. This motivates the following fundamental question:
@@
Can we design an efficient and approximately optimal MARL algorithm for policy-learning in a cooperative multi-agent system with many agents?
Contributions.
We answer this question affirmatively. Our key contributions are outlined below.
@@
  • Subsampling Algorithm. We propose SUBSAMPLE-MFQ, an algorithm designed to address the challenge of MARL with a large number of local agents. We model the problem as a Markov Decision Process (MDP) with a global agent and $n$ local agents. SUBSAMPLE-MFQ selects $k\leq n$ local agents to learn a deterministic policy $\widehat{\pi}^{\mathrm{est}}_{k}$, by applying mean-field value iteration on a subsystem with $k$ local agents to learn $\widehat{Q}_{k}^{\mathrm{est}}$, which can be viewed as a smaller $Q$ function. It then deploys a stochastic policy $\widehat{\pi}_{k}$ which works as follows: the global agent uniformly samples $k$ local agents at each step and uses $\widehat{\pi}_{k}$ to determine its action, whereas each local agent uniformly samples $k-1$ other local agents and uses $\widehat{\pi}_{k}$ to determine its action.
@@
  • Sample Complexity and Theoretical Guarantee. As the number of local agents increases, the size of $\widehat{Q}_{k}$ scales polynomially with $k$, rather than polynomially with $n$ as in mean-field MARL. Analogously, when the size of the local agent’s state space grows, the size of $\widehat{Q}_{k}$ scales exponentially with $k$, rather than exponentially with $n$, as in traditional $Q$-learning). The key analysis technique underlying our results is a novel MDP sampling result. Through it, we show that the performance gap between ${{\pi}_{k}^{\mathrm{est}}}$ and the optimal policy ${\pi^{*}}$ is $\widetilde{O}(1/\sqrt{k})$. The choice of $k$ reveals a fundamental trade-off between the size of the $Q$-table and the optimality of ${\pi}_{k}^{\mathrm{est}}$. When $k\!=\!O(\log n)$, SUBSAMPLE-MFQ is the first MARL algorithm to achieve a run-time that is polylogarithmic in $n$, representing an exponential speedup over the previously best-known polytime mean-field MARL methods, while maintaining a decaying optimality gap.
  • Numerical Simulations. We evaluate the effectiveness of SUBSAMPLE-MFQ in two scenarios: a bounding box problem, and a Gaussian squeeze problem. Our experiments reveal a monotonic improvement in the learned policies as $k\to n$, providing a substantial speedup over mean-field $Q$-learning.
@@
While our results are theoretical in nature, it is our hope that SUBSAMPLE-MFQ will lead to further exploration into the potential of subsampling in general stochastic/Markovian games and networked multi-agent reinforcement learning, and inspire the development of practical algorithms for multi-agent settings.
1.1 Related Literature
@@
MARL has a rich history, starting with early works on Markov games used to characterize the decision-making process (Littman, 1994; Sutton et al., 1999), which can be regarded as a multi-agent extension of the Markov Decision Process (MDP). MARL has since been actively studied (Zhang et al., 2021) in a broad range of settings. MARL is most similar to the category of “succinctly described” MDPs (Blondel & Tsitsiklis, 2000), where the state/action space is a product space formed by the individual state/action spaces of multiple agents, and where the agents interact to maximize an objective. A promising line of research that has emerged over recent years constrains the problem to sparse networked instances to enforce local interactions between agents (Qu et al., 2020a; Lin et al., 2020; Mondal et al., 2022). In this formulation, the agents correspond to vertices on a graph who only interact with nearby agents. By exploiting Gamarnik’s correlation decay property from combinatorial optimization (Gamarnik et al., 2009), they overcome the curse of dimensionality by simplifying the problem to only search over the policy space derived from the truncated graph to learn approximately optimal solutions. However, as the underlying network structure becomes dense with many local interactions, the neighborhood of each agent gets large, causing these algorithms to become intractable.
@@
Mean-Field RL. Under assumptions of homogeneity in the state/action spaces of the agents, the problem of densely networked multi-agent RL was answered in Yang et al. (2018); Gu et al. (2021) by approximating the solution with a mean-field approach where the approximation error scales in $O(1/\sqrt{n})$. To avoid designing algorithms on probability spaces, they study MARL under Pareto optimality and use the law of large numbers to consider a lifted space with a mean agent that aggregates the system’s rewards and dynamics. It then applies kernel regression on $\epsilon$-nets of the lifted space to design policies in time polynomial in $n$. In contrast, our work achieves subpolynomial runtimes by directly sampling from this mean-field distribution. Cui & Koeppl (2022) introduce heterogeneity to mean-field MARL by modeling non-uniform interactions through graphons; however, these methods make critical assumptions on the existence of a sequence of graphons converging in cut-norm to the finite instance. In the cooperative setting, Cui et al. (2023) considers a mean-field setting with $q$ types of homogeneous agents; however, their approach does not converge to the optimum policy.
Structured RL. Our work is related to factored MDPs and exogenous MDPs. In factored MDPs, there is a global action affecting every agent whereas in our case, each agent has its own action (Min et al., 2023; Lauer & Riedmiller, 2000). Our result has a similar flavor to MDPs with exogenous inputs from learning theory, (Dietterich et al., 2018; Foster et al., 2022; Anand & Qu, 2024), where our subsampling algorithm treats each sampled state as an endogenous state, but where the exogenous dependencies can be dynamic.
@@
Miscellaneous. Our work adds to the growing literature on the Centralized Training with Decentralized Execution paradigm (Wang et al., 2023; Zhou et al., 2023), as our algorithm learns a provably approximately optimal policy using centralized information, but makes decisions using local information during execution. In the distributed setting, recent works on V-learning (Jin et al., 2020) reduce the dependence of the product action space to an additive dependence. In contrast, our work further accomplishes a reduction on the complexity of the joint state space, which has not been previously accomplished for dense graphs. Finally, one can approximate the large $Q$-table through linear function approximation (Jin et al., 2020) which significantly reduces the computational complexity. However, achieving theoretical bounds on the performance loss caused by the function approximation is generally intractable without making stronger assumptions such as Linear Bellman completeness (Golowich & Moitra, 2024) or low Bellman-Eluder dimension (Jin et al., 2021a). While our work primarily consider a tabular setting with finite spaces, we provide a further extension of it to the non-tabular setting, under assumptions of Linear Bellman completeness.
@@ 2 Preliminaries
In this section, we formally introduce the problem, state some examples for our setting, and provide technical details of the mean-field and $Q$-learning techniques that will be used throughout the paper.
@@
For $k,n\in\mathbb{N}$ where $k\leq n$, let $\binom{[n]}{k}$ denote the set of $k$-sized subsets of $[n]=\{1,\dots,n\}$. For any vector $z\in\mathbb{R}^{d}$, let $\|z\|_{1}$ and $\|z\|_{\infty}$ denote the standard $\ell_{1}$ and $\ell_{\infty}$ norms of $z$ respectively. Let $\|\mathbf{A}\|_{1}$ denote the matrix $\ell_{1}$-norm of $\mathbf{A}\in\mathbb{R}^{n\times m}$. Given a collection of variables $s_{1},\dots,s_{n}$ the shorthand $s_{\Delta}$ denotes the set $\{s_{i}:i\in\Delta\}$ for $\Delta\subseteq[n]$. We use $\widetilde{O}(\cdot)$ to suppress polylogarithmic factors in all problem parameters except $n$. For a discrete measurable space $(\mathcal{X},\mathcal{F})$, the total variation distance between probability measures $\rho_{1},\rho_{2}$ is given by $\mathrm{TV}(\rho_{1},\rho_{2})=\frac{1}{2}\sum_{x\in\mathcal{X}}|\rho_{1}(x)-\rho_{2}(x)|$. Next, $x\sim\mathcal{D}(\cdot)$ denotes that $x$ is a random element sampled from a probability distribution $\mathcal{D}$, and we denote the uniform distribution over a finite set $\Omega$ by $\mathcal{U}(\Omega)$.
@@ 2.1 Problem Formulation
We consider a system of $n+1$ agents, where agent $g$ is a “global decision making agent” and the remaining $n$ agents, denoted by $[n]$, are “local agents.” At time step $t$, the agents are in state $s(t)=(s_{g}(t),s_{1}(t),...,s_{n}(t))\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, where $s_{g}(t)\in\mathcal{S}_{g}$ denotes the global agent’s state, and for each $i\in[n]$, $s_{i}(t)\in\mathcal{S}_{l}$ denotes the state of the $i$’th local agent. The agents cooperatively select actions $a(t)=(a_{g}(t),a_{1}(t),...,a_{n}(t))\in\mathcal{A}$ where $a_{g}(t)\in\mathcal{A}_{g}$ denotes the global agent’s action and $a_{i}(t)\in\mathcal{A}_{l}$ denotes the $i$’th local agent’s action. At each time-step $t$, the next state for all the agents is independently generated by stochastic transition kernels $P_{g}:\mathcal{S}_{g}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\to[0,1]$ and $P_{l}:\mathcal{S}_{l}\times\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\to[0,1]$ as follows:
$s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),$ (1)
$s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t)),\forall i\in[n].$ (2)
@@
The system then collects a structured stage reward $r(s(t),a(t))$ where the reward $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ depends on $s(t)$ and $a(t)$ through Eq. 3, and where the choice of functions $r_{g}$ and $r_{l}$ is typically application specific.
$r(s,a)=\underbrace{r_{g}(s_{g},a_{g})}_{{\text{global component}}}+\frac{1}{n}\sum_{i\in[n]}\underbrace{r_{l}(s_{i},s_{g},a_{i})}_{\text{local component}},\qquad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (3)
We define a policy $\pi$ as a mapping from $\mathcal{S}$ to $\mathcal{A}$ where we want the policy $\pi$ to maximize the value function for each $s\in S$, which is the expected discounted reward $V^{\pi}(s)=\mathbb{E}_{a(t)\sim\pi(\cdot|s)}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s\right]$, where $\gamma\in(0,1)$ is a discounting factor.
Notably, the cardinality of the search space simplex for the optimal policy is $|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n}$, which is exponential in the number of agents. When noting that the local agents are all homogeneous, and therefore permutation-invariant with respect to the rewards of the system (the order of the other agents does not matter to any single decision-making agent), techniques from mean-field MARL restrict the cardinality of the search space simplex for the optimal policy to $|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|n^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}$, reducing the exponential complexity on $n$ to a polynomial complexity on $n$. In practical systems, when $n$ is large, the $\mathrm{poly}(n)$ sample-complexity may still be computationally infeasible. Therefore, the goal of this problem is to learn an approximately optimal policy with subpolynomial sample complexity, further overcoming the curse of dimensionality.
Example 2.1 (Gaussian Squeeze).
In this task, $n$ homogeneous agents determine their individual action $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, $a_{i}=\{0,\dots,9\}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. In scenarios of traffic congestion, each agent $i\in[n]$ is a traffic controller trying to send $a_{i}$ vehicles into the main road, where controllers coordinate with each other to avoid congestion, hence avoiding either over-use or under-use, thereby contributing to the entire system. This GS problem serves as an ablation study on the impact of subsampling for MARL.
Example 2.2 (Constrained Exploration).
Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent wishes to constrain the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect stage rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, by only using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where some shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Therefore, through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while avoiding unnecessary risk or inefficiency.
Figure 1: Constrained exploration for warehouse accidents.
Figure 2: Traffic congestion settings with Gaussian squeeze.
Due to space constraints, we leave the details of the experiments to Appendix J.
@@
We assume that the state and action spaces of all the agents in the MARL game are finite: $|\mathcal{S}_{l}|,|\mathcal{S}_{g}|,|\mathcal{A}_{g}|,|\mathcal{A}_{l}|<\infty$. Appendix H weakens this assumption to the non-tabular setting of infinite compact sets.
@@
The global and local components of the reward function are bounded. Specifically, $\|r_{g}(\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{g}$, and $\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{l}$. This implies that $\|r(\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{g}+\widetilde{r}_{l}:=\widetilde{r}$.

### v3

@@ Abstract
Designing efficient algorithms for multi-agent reinforcement learning (MARL) is fundamentally challenging because the size of the joint state and action spaces grows exponentially in the number of agents. These difficulties are exacerbated when balancing sequential global decision-making with local agent interactions. In this work, we propose a new algorithm SUBSAMPLE-MFQ (Subsample-Mean-Field-Q-learning) and a decentralized randomized policy for a system with $n$ agents. For any $k\leq n$, our algorithm learns a policy for the system in time polynomial in $k$. We prove that this learned policy converges to the optimal policy on the order of $\tilde{O}(1/\sqrt{k})$ as the number of subsampled agents $k$ increases. In particular, this bound is independent of the number of agents $n$.
@@ 1 Introduction
Reinforcement Learning (RL) has become a popular framework to solve sequential decision making problems in unknown environments and has achieved tremendous success in a wide array of domains such as playing the game of Go (Silver et al., 2016), robotic control (Kober et al., 2013), and autonomous driving (Kiran et al., 2022; Lin et al., 2023a). A key feature of most real-world systems is their uncertain nature, and thus, RL has emerged as a powerful tool for learning optimal policies for multi-agent systems to operate in unknown environments (Kim and Giannakis, 2017; Zhang et al., 2021; Lin et al., 2024; Anand and Qu, 2024). While early RL works focused on the single-agent setting, multi-agent RL (MARL) has recently achieved impressive success in many applications, such as coordinating robotic swarms (Preiss et al., 2017; DeWeese and Qu, 2024), real-time bidding (Jin et al., 2018), ride-sharing (Li et al., 2019), and stochastic games (Jin et al., 2020).
@@
Despite growing interest in MARL, extending RL to multi-agent settings poses significant computational challenges due to the curse of dimensionality. MARL is fundamentally difficult as agents in the real-world not only interact with the environment but also with each other (Shapley, 1953): if each of the $n$ agents has a state space $S$ and action space $A$, the global state-action space has size $(|S||A|)^{n}$, which is exponential in $n$. Thus, many RL algorithms (such as temporal difference learning and tabular $Q$-learning) require computing and storing an $(|S||A|)^{n}$-sized Q-table (Sutton et al., 1999; Bertsekas and Tsitsiklis, 1996). This scalability issue has been observed in a variety of MARL settings (Blondel and Tsitsiklis, 2000; Papadimitriou and Tsitsiklis, 1999; Littman, 1994).
@@
An exciting line of work that addresses this intractability is mean-field MARL (Lasry and Lions, 2007; Yang et al., 2018; Gu et al., 2021, 2022; Hu et al., 2023). The mean-field approach assumes agents are homogeneous in their state-action spaces, enabling their interactions to be approximated by a two-agent setting: here, each agent interacts with a representative “mean agent” which evolves as the empirical distribution of states of all other agents. With these assumptions, mean-field MARL learns an optimal policy with sample complexity $O(n^{|S||A|}|S||A|)$, which is polynomial in the number of agents. However, if $n$ is large, this remains prohibitive even for moderate values of $|S|$ and $|A|$.
@@
Motivated by this problem, in this paper we study the cooperative setting–where agents work collaboratively to maximize a structured global reward–and ask: Can we design a scalable MARL algorithm for learning an approximately optimal policy in a cooperative multi-agent system?
Contributions. We answer this question affirmatively. Our key contributions are outlined below.
@@
Subsampling algorithm. We model the problem as a Markov Decision Process (MDP) with a global agent and $n$ local agents. We propose SUBSAMPLE-MFQ to address the challenge of MARL with a large number of local agents. SUBSAMPLE-MFQ selects $k\leq n$ local agents to learn a deterministic policy $\hat{\pi}^{\mathrm{est}}_{k}$ by applying mean-field value iteration on the $k$-local-agent subsystem to learn $\hat{Q}_{k}^{\mathrm{est}}$, which can be viewed as a smaller $Q$ function. SUBSAMPLE-MFQ then deploys a stochastic policy ${\pi}^{\mathrm{est}}_{k}$, where the global agent samples $k$ local agents uniformly at each step and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action, while each local agent samples $k-1$ other local agents and uses $\hat{\pi}^{\mathrm{est}}_{k}$ to determine its action.
@@
Sample complexity and theoretical guarantee. As the number of local agents increases, the size of $\hat{Q}_{k}$ scales polynomially with $k$, rather than polynomially with $n$ as in mean-field MARL. Analogously, when the size of the local agent’s state space grows, the size of $\hat{Q}_{k}$ scales exponentially with $k$, rather than exponentially with $n$, as in traditional $Q$-learning. The key analytic technique underlying our results is a novel MDP sampling result. Through it, we show that the performance gap between ${{\pi}_{k}^{\mathrm{est}}}$ and the optimal policy ${\pi^{*}}$ is at most $\tilde{O}(1/\sqrt{k})$, with high probability. The choice of $k$ reveals a fundamental trade-off between the size of the $Q$-table and the optimality of ${\pi}_{k}^{\mathrm{est}}$. For example, if $k$ is set to $O(\log n)$, SUBSAMPLE-MFQ is the first centralized MARL algorithm to achieve a polylogarithmic run-time in $n$, representing an exponential speedup over the previously best-known polytime mean-field MARL methods, while maintaining a decaying optimality gap as $n$ gets large,
@@
While our results are theoretical in nature, we hope SUBSAMPLE-MFQ will further exploration of sampling in Markov games, and inspire new practical multi-agent algorithms.
@@
Related work. MARL has a rich history, starting with early works on Markov games (Littman, 1994; Sutton et al., 1999), which are a multi-agent extension of MDPs. MARL has since been actively studied (Zhang et al., 2021) in a broad range of settings. MARL is most similar to the category of “succinctly described” MDPs (Blondel and Tsitsiklis, 2000), where the state/action space is a product space formed by the individual state/action spaces of multiple agents, and where the agents interact to maximize an objective. A recent line of research constrains the problem to sparse networked instances to enforce local interactions between agents (Qu et al., 2020a; Lin et al., 2020; Mondal et al., 2022). In this formulation, the agents correspond to vertices on a graph who only interact with nearby agents. By exploiting Gamarnik’s correlation decay property from combinatorial optimization (Gamarnik et al., 2009), they overcome the curse of dimensionality by simplifying the problem to only search over the policy space derived from the truncated graph to learn approximately optimal solutions. However, as the underlying network structure becomes dense with many local interactions, the neighborhood of each agent gets large, and these algorithms become intractable.
@@
Mean-Field RL. Under assumptions of homogeneity in the state/action spaces of the agents, the problem of densely networked multi-agent RL was studied by Yang et al. (2018); Gu et al. (2021), who approximated the solution in polynomial time with a mean-field approach where the approximation error scales in $O(1/\sqrt{n})$. In contrast, our work achieves subpolynomial runtimes by directly sampling from this mean-field distribution. Cui and Koeppl (2022) introduce heterogeneity to mean-field MARL by modeling non-uniform interactions through graphons; however, these methods crucially assume the existence of graphon sequences that converge in cut-norm to the finite graph. In the cooperative setting, Subramanian et al. (2022); Cui et al. (2023) studies a mean-field setting with $q$ types of homogeneous agents; however, their learned policy does not provably converge to the optimum.
@@
Other related works. Our work is related to factored MDPs, where there is a global action affecting every agent; however, in our case, each agent has its own action (Min et al., 2023; Lauer and Riedmiller, 2000). Jin et al. (2020) reduces the dependence of the product action space to an additive dependence with V-learning. Our work further reduces the complexity of the joint state space, which has not been previously accomplished. We add to the growing literature on the Centralized Training with Decentralized Execution regime (Zhou et al., 2023), as our algorithm learns a provably approximately optimal policy using centralized information, but makes decisions using only local information during execution. Finally, one can efficiently approximate the $Q$-table through function approximation (Jin et al., 2021). However, achieving theoretical bounds on the performance loss due to function approximation is intractable without strong assumptions such as linear Bellman completeness or low Bellman-Eluder dimension (Golowich and Moitra, 2024). While our work primarily studies the finite tabular setting, we extend it to non-tabular linear MDPs in Appendix J.
@@ 2 Preliminaries
For $k,n\in\mathbb{N}$ where $k\leq n$, let $\binom{[n]}{k}$ denote the set of $k$-sized subsets of $[n]=\{1,\dots,n\}$. For any vector $z\in\mathbb{R}^{d}$, let $\|z\|_{1}$ and $\|z\|_{\infty}$ denote the standard $\ell_{1}$ and $\ell_{\infty}$ norms of $z$ respectively. Let $\|\mathbf{A}\|_{1}$ denote the matrix $\ell_{1}$-norm of $\mathbf{A}\in\mathbb{R}^{n\times m}$. Given variables $s_{1},\dots,s_{n}$, $s_{\Delta}\coloneqq\{s_{i}:i\in\Delta\}$ for $\Delta\subseteq[n]$. We use $\tilde{O}(\cdot)$ to suppress polylogarithmic factors in all problem parameters except $n$. For a discrete measurable space $(\mathcal{X},\mathcal{F})$, the total variation distance between probability measures $\rho_{1},\rho_{2}$ is given by $\mathrm{TV}(\rho_{1},\rho_{2})=\frac{1}{2}\sum_{x\in\mathcal{X}}|\rho_{1}(x)-\rho_{2}(x)|$. Next, $x\sim\mathcal{D}(\cdot)$ denotes that $x$ is a random element sampled from a distribution $\mathcal{D}$, and we denote that $x$ is a random sample from the uniform distribution over a finite set $\Omega$ by $x\sim\mathcal{U}(\Omega)$. We include a detailed notation table in Table 1.
@@ 2.1 Problem formulation
We consider a system of $n+1$ agents, where agent $g$ is a “global decision making agent” and the remaining $n$ agents, denoted by $[n]$, are “local agents.” At time $t$, the agents are in state $s(t)=(s_{g}(t),s_{1}(t),...,s_{n}(t))\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, where $s_{g}(t)\in\mathcal{S}_{g}$ denotes the global agent’s state, and for each $i\in[n]$, $s_{i}(t)\in\mathcal{S}_{l}$ denotes the state of the $i$’th local agent. The agents cooperatively select actions $a(t)=(a_{g}(t),a_{1}(t),...,a_{n}(t))\in\mathcal{A}\coloneqq\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$, where $a_{g}(t)\in\mathcal{A}_{g}$ denotes the global agent’s action and $a_{i}(t)\in\mathcal{A}_{l}$ denotes the $i$’th local agent’s action. At time $t+1$, the next state for all the agents is independently generated by stochastic transition kernels $P_{g}:\mathcal{S}_{g}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\to\Delta(\mathcal{S}_{g})$ and $P_{l}:\mathcal{S}_{l}\times\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}\to\Delta(\mathcal{S}_{l})$ as follows:
$s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),\quad s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t)),\forall i\in[n].$ (1)
@@
The system then collects a structured stage reward $r(s(t),a(t))$ where the reward $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ depends on $s(t)$ and $a(t)$ through Equation 2, and where $r_{g}$ and $r_{l}$ is typically application specific.
$r(s,a)=\underbrace{r_{g}(s_{g},a_{g})}_{{\text{global component}}}+\frac{1}{n}\sum_{i\in[n]}\underbrace{r_{l}(s_{i},s_{g},a_{i})}_{\text{local component}}$ (2)
A policy $\pi:\mathcal{S}\to\mathcal{P}(\mathcal{A})$ maps from states to distributions of actions such that $a\sim\pi(\cdot|s)$. Given $\gamma\in(0,1)$, we seek to learn a policy $\pi$ that maximizes the ($\gamma$-discounted) value for each $s\in S$:
$V^{\pi}(s)=\mathbb{E}_{a(t)\sim\pi(\cdot|s)}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s\Big{]}.$ (3)
The cardinality of the search space simplex for the optimal policy is $|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n}$, which is exponential in the number of agents, underscoring the need for efficient approximation algorithms.
@@
We assume that the state and action spaces of all the agents in the MARL game are finite: $|\mathcal{S}_{l}|,|\mathcal{S}_{g}|,|\mathcal{A}_{g}|,|\mathcal{A}_{l}|<\infty$. Appendix J of the supplementary material relaxes this assumption to the non-tabular setting with infinite continuous sets.
@@
The components of the reward function are bounded. Specifically, $\|r_{g}(\cdot,\cdot)\|_{\infty}\leq\tilde{r}_{g}$, and $\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\leq\tilde{r}_{l}$. This implies $\|r(\cdot,\cdot)\|_{\infty}\leq\tilde{r}_{g}+\tilde{r}_{l}:=\tilde{r}$.